# Staircase Cascaded Fusion of Lightweight Local Pattern Recognition and Long-Range Dependencies for Structural Crack Segmentation

### v3

@@ Abstract
Detecting cracks with pixel-level precision for key structures is a significant challenge, existing methods struggle to integrate local textures and pixel dependencies of cracks. Furthermore, these methods possess numerous parameters and substantial computational requirements, complicating deployment on edge devices. In this paper, we propose the Staircase Cascaded Fusion Crack Segmentation Network (CrackSCF), which generates high-quality crack segmentation maps while reducing computational overhead. We design a lightweight convolutional block that substitutes all convolution operations, reducing the model’s computational demands while maintaining an effective capture of local details. Additionally, we introduce a lightweight long-range dependency extractor to better capture the long-range dependencies. Furthermore, we develop a staircase cascaded fusion module, which seamlessly integrates local patterns and long-range dependencies, resulting in high-quality segmentation maps. To comprehensively evaluate our method, we created the challenging TUT benchmark dataset and evaluated it alongside five other public datasets. The results show that our method outperforms existing methods, particularly in handling background noise and achieving detailed segmentation. The F1 and mIoU scores on the TUT dataset are 0.8382 and 0.8473, respectively, demonstrating state-of-the-art (SOTA) performance with low computational resources. The code and dataset is available at https://github.com/Karl1109/CrackSCF.
Index Terms:
@@
Structural crack segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
publicationid: pubid: $0$–$00$ 2025 IEEE
I Introduction
@@
Cracks are common defects in structures such as pavements and buildings, which can lead to serious safety issues. Thus, automated detection on edge control devices is crucial for ensuring production safety [1]. Early automated detection methods used traditional image processing techniques for crack extraction [2, 3, 4]. While these methods are easy to implement, their performance in complex scenarios still needs improvement.
Figure 1: Performance comparison of our proposed method with UCTNet [5], SFIAN [6], CT-crackseg [7], DTrCNet [8], Crackmer [9] and Simcrack [10] on the TUT dataset in terms of mIoUs, FLOPs and Params metrics.
@@
In recent years, benefiting from the strong local inductive capability of Convolutional Neural Networks (CNN), CNN-based networks such as MST-Net [11] and FcaHRNet [12] have shown efficient feature extraction capabilities in structural crack segmentation tasks. The spatial invariance of CNNs allows them to detect cracks anywhere in an image by learning similar features at different locations. In labeled images, crack labels are long, narrow structures of white pixels, while the background is large black areas. This challenges the model to distinguish between thin cracks and the background. Nevertheless, networks based on CNNs face challenges in capturing long-range dependencies between crack pixels, leading to segmentation discontinuities.
@@
The success of Vision Transformer (ViT) [13] and Swin Transformer [14] has demonstrated the advantages of Transformer [15], with their strong sequence processing capabilities, in handling visual tasks [16, 17]. Unlike CNNs, transformers can capture dependencies between any pixels when processing images [18, 19], making them more advantageous in handling complex topologies. To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as CATransUNet [20], MFAFNet [21], and Crackmer [8], which integrate local and long-range information through U-shaped connections or residual connections.
@@
    The different information fusion strategies significantly affect the segmentation performance of CNN-Transformer networks. For instance, UCTNet [5] and CT-crackseg [7] use channel concatenation to fuse local patterns and long-range dependences information. However, concatenation merely stacks the feature information together without deeper feature interaction, failing to understand their correlation and complementarity, resulting in poor segmentation performance on fine cracks and susceptibility to background noise. Additionally, DTrCNet [8] and Crackmer [9] use channel attention to fuse information, which can enhance focus on certain features but neglects spatial information and pixel details in the feature maps, impacting the model’s overall performance. As shown in Fig. 1(a), due to these shortcomings, these four methods fail to achieve the best performance.
@@
    Most networks, in pursuit of performance, overlook the significant parameter and computational costs brought by numerous convolution operations, this makes them challenging to deploy on resource-constrained control devices. As shown in Fig. 1(b) and Fig. 1(c), networks like SFIAN [6], Crackmer [9], and CT-crackseg[7] have very high computational FLOPs and parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
@@
To comprehensively evaluate the structural crack segmentation performance of the model in various complex environments, we also collected a dataset named TUT, which contains 1408 RGB images. Existing public datasets like DeepCrack [22] and Crack500 [23] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance. The TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines. With such a diverse set of image scenarios, this dataset effectively evaluates the network’s generalization and robustness. Furthermore, the images feature occlusions, highly complex backgrounds, and uneven lighting, which further enhances the datasets diversity.
@@
The organization of the subsequent sections of this paper is as follows: In Section II, existing crack detection models and lightweight networks are reviewed. Section III details the components of the proposed network. Section IV delves into the experimental details and results. Finally, Section V provides a summary of the entire paper and highlights future research directions.
II Related Work
@@
II-A Crack Segmentation Networks
@@
Recently, a growing number of researchers have utilized CNNs to develop crack segmentation networks, achieving advanced performance. Al-Huda et al. [24] proposed the MDAUNet network, based on U-Net and dual attention modules, and designed a hybrid weighted segmentation loss function to address class imbalance issues. Cheng et al. [6] introduced the Selective Feature Fusion and Irregular-Aware Network (SFIAN) to effectively model irregular crack objects, resulting in clearer backgrounds on various crack images. Huang et al. [25] developed a crack detection network for masonry structures that combines thermal imaging and visible light image fusion with semantic segmentation, laying the foundation for the development of cross-modal methods for structural crack monitoring.
Although the aforementioned methods have demonstrated the advantages of CNNs in crack segmentation tasks, CNNs are essentially local operations and cannot effectively perceive a small number of pixels amidst a large background. Therefore, these methods often result in discontinuous crack pixels and are highly susceptible to background noise interference.
@@
Benefit from the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes. Therefore, many researchers are amalgamating the outputs of CNNs and transformers for crack segmentation. Chen et al. [26] proposed TransUNet, a novel network that effectively integrates long-range information and local features through a U-Net-like structure. Qi et al. [27] developed an end-to-end model based on ViT and level set theory for segmenting defects in bridge pavement, achieving precise segmentation by amalgamating the outputs of two parallel decoders. Feng et al. [28] developed a shale crack segmentation network that combines CNN and Swin Transformer, utilizing Fast Fourier Transform (FFT) to merge feature maps from both, effectively suppressing high-frequency noise in the imaging process.
The aforementioned methods, while demonstrating the advantages of CNN-Transformer networks in processing pixel sequences and suppressing background noise, have limitations due to using only one method or branch for feature fusion. This reduces their ability to effectively perceive long-range dependencies and local patterns, causing some complex regions’ local or long-range information to be overshadowed. Additionally, these methods excessively focus on performance, overlooking the substantial computational resources required for feature extraction and fusion, making them difficult to deploy on edge computing devices.
II-B Lightweight Networks
@@
Deploying deep learning models on edge devices like smartphones and drones is trending. However, these devices have limited computing resources, making it challenging to use models with high parameter counts. Thus, compact model design for lightweight modifications in size and speed is crucial [29, 30]. For example, Howard et al. [31] proposed MobileNet, an efficient model using depthwise separable convolutions to reduce computational load and parameter count, suitable for resource-constrained devices. Ma et al. [32] introduced ShuffleNet, providing practical design principles for efficient network architecture. Additionally, low-rank approximation compression, effective in fields like numerical linear algebra and image processing, is now used to compress deep learning network architectures. Swaminathan et al. [33] introduced SLR, a sparse low-rank method that improves compression rates with sparse SVD matrices.
@@
In recent years, there has been a growing trend among researchers to integrate model lightweighting techniques with crack segmentation models, aiming to facilitate real-time crack detection on edge computing devices. For example, Zhu et al. [34] designed RHACrackNet, a lightweight crack detection network integrating hybrid attention and residual blocks, enabling real-time detection on edge devices. Huang et al. [35] proposed a lightweight fusion attention module for road crack detection, ensuring high-precision crack extraction with rapid detection. Although these lightweight models reduce computational resources, their emphasis on minimizing parameters results in limited model expressiveness, hindering their ability to effectively capture intricate patterns and features in the input data. This leads to unsatisfactory segmentation results. This imbalance between performance and lightweight design makes them unreliable in practical applications.
III Methodology
III-A Network Architecture
@@
Specifically, for a given dataset, the crack images are first input into the MFE module, which is divided into four stages. In each stage, the extracted feature maps undergo a Feature Enhancement stage. Then, the LDE module processes these enhanced feature maps to obtain pixel sequences rich in long-range dependencies. Subsequently, the SCFM module receives these sequences along with the four layers of feature maps extracted by the MFE. After processing through four stages, the resolution is doubled at each stage while the number of channels is halved, ultimately resulting in the segmentation result.
@@
Figure 3: Comparison of FLOPs and Params for each module using the original convolution and our proposed LRDS convolution block.
III-B Lightweight Multi-scale Feature Extractor
As shown in Fig. 4, our designed LRDS block adopts a bottleneck-like structure. It comprises a Reduce layer, a depthwise convolution layer, a pointwise convolution layer, and an Expand layer. The input feature map first passes through the Reduce layer, reducing the input channels to $C_{m}$. It then goes through the depthwise convolution layer and the pointwise convolution layer, keeping the number of channels unchanged. Finally, the Expand layer increases the number of channels to $C_{d}$, producing the final output feature map.
@@ III-B Lightweight Multi-scale Feature Extractor
Low-rank approximation theory greatly reduces the convolutional layer parameters and computational load. In order to efficiently capture the spatial information of feature maps in low-rank space for finer perception of detailed regions, we process the features in LRDS by embedding deep convolution and pointwise convolution in low-rank space. We constructed MFE by replacing all convolution operations in ResNet50 with LRDS convolution blocks.
Figure 4: Illustration of LRDS convolution block. After the number of channels of the input features is reduced, their local and spatial features are efficiently extracted by depth-wise convolution and point-wise convolution, and finally the number of channels is restored.
III-C Lightweight Long-range Dependency Extractor
@@
$\centering\begin{split}&\text{MSDeformAttn}(z_{q},\hat{p}_{q},\{x_{c}\}_{c=1}^{C})=\\ &\sum_{h=1}^{H}W_{h}\Bigg{[}\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}\cdot W^{\prime}_{h}x_{c}(\phi_{c}(\hat{p}_{q})+\Delta p_{hcqt})\Bigg{]}\end{split}\@add@centering$ (4)
@@
$M=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Z)))$ (9)
@@
$Y=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Y_{0})))$ (11)
@@
III-D Staircase Cascaded Fusion Module
Figure 6: Illustration of the PAF. It effectively reduces required computing resources by embedding LRDS.
@@ III-D Staircase Cascaded Fusion Module
$\sigma=\text{Sigmoid}(f_{c}(\vec{v_{c}})\cdot f_{d}(\vec{v_{d}}))$ (12)
where $f_{c}$ and $f_{d}$ denote pointwise convolution and batch normalization operations, respectively. Thus, the output of PAF can be represented as:
$\text{Output}=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (13)
@@
$\begin{split}&F_{i}^{\text{O}}=C_{\text{i}}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\right)\odot P_{i}\left(F_{i}^{\text{MFE}},F_{i-1}^{\text{O}}\right)\end{split}$ (14)
@@
III-E Loss Function
The discrepancy between predictions and actual values is quantified by the loss function; a smaller value for this loss suggests superior model performance. Previously, Binary Cross-Entropy (BCE) [36] was often used, but it may not effectively distinguish the minority crack pixels in the image. Combining BCE with Dice loss [37] addresses this issue. BCE improves the probability distribution for each pixel, while Dice loss maximizes the overlap between predictions and true labels, enhancing robustness against imbalanced data.
@@ IV Experiments
IV-A Dataset Description
TABLE I: Comparison of Different Crack Datasets
@@ IV-A Dataset Description
6 AEL [23] 2019 58 768 width and 512 height 0.67% bitumen
7 CrackTree206 [41] 2012 206 800 width and 600 height 0.32% bitumen
8 LSCD [42] 2024 143 400 width and 400 height 3.18% bitumen, cement
9 CrackSC [43] 2023 197 320 width and 480 height 2.02% bitumen, cement
@@
Figure 7: Illustration of the images of each scenarios in the TUT dataset.
The previous works has used dataset scenarios that are relatively simple, such as cracks in cement or bitumen pavements only. These datasets have clear backgrounds and minimal interference, which cannot comprehensively evaluate the performance of models. To tackle this problem, we created the TUT dataset to enable better segmentation across different scenarios. Detailed information and comparisons with other datasets are list in Table I.
The TUT dataset comprises 1270 crack images captured with our mobile phones, along with a small number of 138 images sourced from the Internet. TUT dataset contains 900 images of bitumen and cement with irregular crack morphology but flat backgrounds, 142 images of bricks and tiles with regular crack morphology, 228 images of plastic runways with extremely complex cracks, and a total of 138 images of metals, generator blades, and underground pipelines with complex lighting and background disturbances. To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images. Unlike other datasets with clear backgrounds, the TUT dataset images have complex, noisy backgrounds and cracks of intricate shapes. This evaluates the model’s capability to segment cracks in challenging conditions.
As list in Table I, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios. Fig. 7 shows crack images in different scenarios and their binary labeled images. The crack pixel ratio in the TUT dataset is 3.16%, which is moderate. This ensures the model learns useful information without being hindered by too few crack pixels and prevents the model from overly relying on the crack pixel ratio when there are too many crack pixels.

### v4

# CrackSCF: Lightweight Cascaded Fusion Network for Robust and Efficient Structural Crack Segmentation

@@ Abstract
Accurately segmenting structural cracks at the pixel level remains a major hurdle, as existing methods fail to integrate local textures with pixel dependencies, often leading to fragmented and incomplete predictions. Moreover, their high parameter counts and substantial computational demands hinder practical deployment on resource-constrained edge devices. To address these challenges, we propose CrackSCF, a Lightweight Cascaded Fusion Crack Segmentation Network designed to achieve robust crack segmentation with exceptional computational efficiency. We design a lightweight convolutional block (LRDS) to replace all standard convolutions. This approach efficiently captures local patterns while operating with a minimal computational footprint. For a holistic perception of crack structures, a lightweight Long-range Dependency Extractor (LDE) captures global dependencies. These are then intelligently unified with local patterns by our Staircase Cascaded Fusion Module (SCFM), ensuring the final segmentation maps are both seamless in continuity and rich in fine-grained detail. To comprehensively evaluate our method, we created the challenging TUT benchmark dataset and evaluated it alongside five other public datasets. The experimental results show that the CrackSCF method consistently outperforms the existing methods, and it demonstrates greater robustness in dealing with complex background noise. On the TUT dataset, CrackSCF achieved 0.8382 on F1 score and 0.8473 on mIoU, and it only required 4.79M parameters. The code and dataset is available at https://github.com/Karl1109/CrackSCF.
{IEEEkeywords}
@@
Crack segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
1 Introduction
[…]
@@ 1 Introduction
Cracks are common defects in industrial structures such as pavements and buildings, which can lead to serious safety issues. Thus, automated detection on edge control devices is crucial for ensuring production safety [1]. Early automated detection methods used traditional image processing techniques for crack extraction [2]. While these methods are easy to implement, their performance in complex scenarios still needs improvement.
Figure 1: (a) shows the performance of our method compared to other methods. (b) shows the different scenarios included in TUT dataset.
@@
In recent years, benefiting from the strong local inductive capability of Convolutional Neural Networks (CNN), CNN-based networks such as Crack500 [3] and SC-OSDA [4] have shown efficient feature extraction capabilities in structural crack segmentation tasks. The spatial invariance of CNNs allows them to detect cracks anywhere in an image by learning similar features at different locations. In labeled images, crack labels are long, narrow structures of white pixels, while the background is large black areas. This challenges the model to distinguish between thin cracks and the background. Nevertheless, networks based on CNNs face challenges in capturing long-range dependencies between crack pixels, leading to segmentation discontinuities.
@@
The success of Vision Transformer (ViT) [5] has demonstrated the advantages of Transformer, with their strong sequence processing capabilities, in handling visual tasks [6]. Unlike CNNs, transformers can capture dependencies between any pixels when processing images, making them more advantageous in handling complex topologies. To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as DTrCNet [7] and Crackmer [7], which integrate local and long-range information through U-shaped or residual connections.
@@
    The different fusion strategies significantly affect the segmentation performance of CNN-Transformer networks. For instance, UCTNet [8] and CTCS [9] use channel concatenation to fuse local patterns and long-range dependences information. However, concatenation merely stacks the feature information together without deeper feature interaction, failing to understand their correlation and complementarity, resulting in poor segmentation performance on fine cracks and susceptibility to background noise. Additionally, DTrCNet [7] and Crackmer [10] use channel attention to fuse information, which can enhance focus on certain features but neglects spatial information and pixel details in the feature maps, impacting the model’s overall performance.
@@
    Most networks, in pursuit of performance, overlook the parameter and computational costs brought by numerous convolution operations, this makes them challenging to deploy on resource-constrained control devices. As shown in Fig. 1(a), networks like SFIAN [11], Crackmer [10], and CTCS[9] have high computational parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
@@
To comprehensively evaluate the structural crack segmentation performance of the model in various complex environments, we also collected a dataset named TUT, which contains 1408 RGB images. Existing public datasets like DeepCrack [12] and Crack500 [3] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance. As shown in Fig. 1(b), the TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines. With such a diverse set of image scenarios, this dataset effectively evaluates the network’s generalization and robustness. Furthermore, the images feature occlusions, highly complex backgrounds, and uneven lighting, which further enhances the datasets diversity.
@@
The organization of the subsequent sections of this paper is as follows: In Section 2, existing crack detection models and lightweight networks are reviewed. Section 3 details the components of the proposed network. Section 4 delves into the experimental details and results. Finally, Section 5 provides a summary of the entire paper and highlights future research directions.
2 Related Work
@@
2.1 Crack Segmentation Networks
@@
Recently, a growing number of researchers have utilized CNNs to develop crack segmentation networks, achieving advanced performance. Zhu et al. [13] proposed the BSNBM network, which integrates bidirectional self-rectifying modules with Bayesian inference to enhance crack detection robustness and accuracy under uncertainty. Cheng et al. [11] introduced the Selective Feature Fusion and Irregular-Aware Network (SFIAN) to effectively model irregular crack objects, resulting in clearer backgrounds on various crack images. Lyu et al. [14] proposed a CNN-based crack quantification framework that utilizes voxel-based reconstruction and Bayesian data fusion to achieve accurate crack width estimation.
@@
Benefit from the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes. Therefore, many researchers are amalgamating the outputs of CNNs and transformers for crack segmentation. Qi et al. [15] developed an end-to-end model based on ViT and level set theory for segmenting defects in bridge pavement, achieving precise segmentation by amalgamating the outputs of two parallel decoders. Feng et al. [16] developed a shale crack segmentation network that combines CNN and Swin Transformer, utilizing Fast Fourier Transform (FFT) to merge feature maps from both, effectively suppressing high-frequency noise in the imaging process.
Although CNN and CNN-Transformer methods have shown promise in crack segmentation, CNNs struggle to capture global context due to their local receptive fields, leading to discontinuous predictions and vulnerability to background noise. CNN-Transformer hybrids improve global perception but often adopt single-branch fusion, limiting their ability to integrate local and long-range cues. Moreover, both lack selective interaction and noise suppression, and their high computational cost hinders deployment on edge devices.
2.2 Lightweight Networks
@@
Deploying deep learning models on edge devices like smartphones and drones is trending. However, these devices have limited computing resources, making it challenging to use models with high parameter counts. Thus, compact model design for lightweight modifications in size and speed is crucial. For example, Howard et al. [17] proposed MobileNet, an efficient model using depthwise separable convolutions to reduce computational load and parameter count, suitable for resource-constrained devices. Ma et al. [18] introduced ShuffleNet, providing practical design principles for efficient network architecture. Additionally, low-rank approximation compression, effective in fields like numerical linear algebra and image processing, is now used to compress deep learning network architectures. Swaminathan et al. [19] introduced SLR, a sparse low-rank method that improves compression rates with sparse SVD matrices.
@@
In recent years, there has been a growing trend among researchers to integrate model lightweighting techniques with crack segmentation models, aiming to facilitate real-time crack detection on edge computing devices. For example, Zhu et al. [20] designed RHACrackNet, a lightweight crack detection network integrating hybrid attention and residual blocks, enabling real-time detection on edge devices. Huang et al. [21] proposed a lightweight fusion attention module for road crack detection, ensuring high-precision crack extraction with rapid detection. Although lightweight models reduce computational cost, their aggressive parameter reduction limits model expressiveness, impairing the capture of intricate patterns and leading to suboptimal segmentation. This trade-off between performance and efficiency undermines their reliability in real-world applications.
3 Methodology
3.1 Network Architecture
@@
Specifically, for a given dataset, the images are first input into the MFE, which is divided into four stages. In each stage, the extracted feature maps undergo a Feature Enhancement stage. Then, the LDE processes these enhanced feature maps to obtain pixel sequences rich in long-range dependencies. Subsequently, the SCFM receives these sequences along with the four layers of feature maps extracted by the MFE. After processing through four stages, the resolution is doubled at each stage while the number of channels is halved, ultimately resulting in the segmentation result.
@@
Figure 3: Comparison of each module when using our LRDS or original convolution separately. (a) shows the comparison of FLOPs, (b) shows the comparison of parameters.
3.2 Lightweight Multi-scale Feature Extractor
As shown in Fig. 2, our designed LRDS block adopts a bottleneck-like structure. It comprises a Reduce layer, a depthwise convolution layer, a pointwise convolution layer, and an Expand layer. The input feature map first passes through the Reduce layer, reducing the input channels to $f_{mid}$. It then goes through the depthwise convolution layer and the pointwise convolution layer, keeping the number of channels unchanged. Finally, the Expand layer increases the number of channels to $f_{out}$, producing the final output feature map.
@@ 3.2 Lightweight Multi-scale Feature Extractor
In order to efficiently capture the spatial information of feature maps in low-rank space for finer perception of detailed regions, we process the features in LRDS by embedding deep convolution and pointwise convolution in low-rank space. We constructed MFE by replacing all convolution operations with LRDS convolution blocks.
3.3 Lightweight Long-range Dependency Extractor
@@
$\centering\begin{split}&MSDeformAttn(z_{q},\hat{p}_{q},\{x_{c}\}_{c=1}^{C})=\\ &\sum_{h=1}^{H}W_{h}\Bigg[\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}\cdot W^{\prime}_{h}x_{c}(\phi_{c}(\hat{p}_{q})+\Delta p_{hcqt})\Bigg]\end{split}\@add@centering$ (4)
@@
M=LR_Linear(GELU(LR_Linear(Z)))M=LR\_Linear(GELU(LR\_Linear(Z))) (9)
@@
$Y=LR\_Linear(GELU(LR\_Linear(Y_{0})))$ (11)
@@
3.4 Staircase Cascaded Fusion Module
@@
$Output=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (12)
@@
$\begin{split}&F_{i}^{O}=C_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\right)\odot P_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\end{split}$ (13)
@@
3.5 Loss Function
The discrepancy between predictions and actual values is quantified by the loss function; a smaller value for this loss suggests superior model performance. Previously, Binary Cross-Entropy (BCE) was often used, but it may not effectively distinguish the minority crack pixels in the image. Combining BCE with Dice loss addresses this issue. BCE improves the probability distribution for each pixel, while Dice loss maximizes the overlap between predictions and true labels, enhancing robustness against imbalanced data [22] .
@@ 4 Experiments
4.1 Dataset Description
Table 1: Comparison of Different Crack Datasets
@@ 4.1 Dataset Description
The previous works has used dataset scenarios that are relatively simple, such as cracks in cement or bitumen pavements only. These datasets have clear backgrounds and minimal interference, which cannot comprehensively evaluate the performance of models. To tackle this problem, we created the TUT dataset to enable better segmentation across different scenarios. Detailed information and comparisons with other datasets are list in Table 1.
The TUT dataset comprises 1270 crack images, along with a small number of 138 images sourced from the Internet. TUT dataset contains 900 images of bitumen and cement with irregular crack morphology but flat backgrounds, 142 images of bricks and tiles with regular crack morphology, 228 images of plastic runways with extremely complex cracks, and a total of 138 images of metals, generator blades, and underground pipelines with complex lighting and background disturbances. To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images. Unlike other datasets with clear backgrounds, the TUT dataset images have complex backgrounds and cracks of intricate shapes. This evaluates the model’s capability to segment cracks in challenging conditions.
Figure 5: Illustration of the various scenarios of the TUT dataset with the original images and Ground Truth.
As list in Table 1, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios. Fig. 5 shows crack images in different scenarios and their binary labeled images. The crack pixel ratio in the TUT dataset is 3.16%, which is moderate. This ensures the model learns useful information without being hindered by too few crack pixels and prevents the model from overly relying on the crack pixel ratio when there are too many crack pixels.